OpenStudy (anonymous):

Find the probability of randomly drawing the given marbles from a bag of 4 red, 6 green, and 2 blue marbles (a) with replacement and (b) without replacement. 26. a red, then a green

OpenStudy (across):

There is a total of \(12\) marbles in the bag. The probabilities without replacement are as follows:\[P(\text{red})=1/3,\\P(\text{green})=1/2,\\P(\text{blue})=1/6.\]Therefore, \(P(\text{red and green})=1/6\). Can you do the one without replacement?

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

ok yeah dont know how to do the second part